# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

### v1

@@ Appendix C Network Centrality Measures per Community
Table C presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc. For each community, we report basic network metrics (number of nodes, edges, average degree, weighted degree, clustering coefficient, and average shortest path length) as well as the top 5 most central nodes according to four different centrality measures (betweenness, closeness, degree, and eigenvector centrality). The whole network shows high connectivity (average degree 40.28) and clustering (0.52), with core hashtags like “addiction”, “recovery”, and “sober” consistently ranking highly across all centrality measures.
@@
Table 4: Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.
Network #N #E Avg D Avg WD Avg C Avg SP Betweenness Closeness Degree Eigenvector
Whole Network 2333 46990 40.28 513.84 0.52 2.20 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
Health Conditions 277 1522 10.99 224.09 0.85 2.02 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
Platform 183 146 1.60 39.58 0.12 2.44 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp:) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
Identity & Community 155 97 1.25 8.22 0.09 3.21 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
Location 116 19 0.33 1.26 0.00 1.80 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
Cannabis 51 229 8.98 39.25 0.55 2.04 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
Alcohol 43 92 4.28 56.19 0.41 1.79 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
Substance Effects 41 33 1.61 23.76 0.16 2.42 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
Consumption Method 41 62 3.02 89.80 0.30 1.86 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
[…]
Occupation 73 5 0.14 0.93 0.00 1.50 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
Misc 565 2367 8.38 139.01 0.29 2.46 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
Table 5: Centrality measures for the whole network and each community. #N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path. The top five nodes are presented per centrality measure.
Appendix D Hashtag Community Distribution per Content Topic
Table 4 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content. This analysis reveals notable patterns in how hashtags from different communities are employed across content types. Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities. Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content. Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities. This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.

### v2

@@ Appendix A.2 Network Centrality Measures per Community
Table A.1 presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc. For each community, we report basic network metrics (number of nodes, edges, average degree, weighted degree, clustering coefficient, and average shortest path length) as well as the top 5 most central nodes according to four different centrality measures (betweenness, closeness, degree, and eigenvector centrality). The whole network shows high connectivity (average degree 40.28) and clustering (0.52), with core hashtags like “addiction”, “recovery”, and “sober” consistently ranking highly across all centrality measures.
Network #N #E Avg D Avg WD Avg C Avg SP Avg B Betweenness Closeness Degree Eigenvector
Whole Network 2333 46990 40.28 513.84 0.52 2.20 0.0005 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 0.0010 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
Health Conditions 277 1522 10.99 224.09 0.85 2.02 0.0012 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 0.0002 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
Platform 183 146 1.60 39.58 0.12 2.44 0.0004 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp (shi) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
Identity & Community 155 97 1.25 8.22 0.09 3.21 0.0001 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 0.0007 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
Location 116 19 0.33 1.26 0.00 1.80 ¡0.0001 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
Cannabis 51 229 8.98 39.25 0.55 2.04 0.0006 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
Alcohol 43 92 4.28 56.19 0.41 1.79 0.0006 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 0.0001 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
Substance Effects 41 33 1.61 23.76 0.16 2.42 0.0005 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
Consumption Method 41 62 3.02 89.80 0.30 1.86 0.0017 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 0.0004 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
Other Substances 70 130 3.71 50.11 0.27 2.69 0.0012 narcan, snow, psychedelictok, naloxone, naltrexone narcan, psychedelictok, snow, , shroooms psychedelictok, snow, narcan, mushrooms, naloxone psychedelictok, mushrooms, salvia, mescaline, magicmushroomsadventures
Humor 56 19 0.68 28.64 0.00 1.90 0.0002 addictionhumor, prank, pingtok⚫️ ⚫️, standupcomedy, satire addictionhumor, jokes, comedia, soberhumor, comediahumor addictionhumor, jokes, comedia, soberhumor, comediahumor addictionhumor, jokes, comedia, soberhumor, comediahumor
Occupation 73 5 0.14 0.93 0.00 1.50 ¡0.0001 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
Misc 565 2367 8.38 139.01 0.29 2.46 0.0002 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
Table A.1. Centrality measures for the whole network and each community. #N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path, Avg B=Average Betweenness Centrality. The top five nodes are presented per centrality measure. Consumption Method exhibits the highest average betweenness (bolded).
Appendix A.3 Hashtag Community Distribution per Content Topic
Table A.2 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content. This analysis reveals notable patterns in how hashtags from different communities are employed across content types. Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities. Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content. Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities. This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.
@@ Appendix A.3 Hashtag Community Distribution per Content Topic
Table A.2. Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.